Orchid Biotech Company is evaluating several different development projects for experimental drugs. Although the cash flows are difficult to​ forecast, the company has come up with the following estimates of the initial capital requirements and NPVs for the projects. Given a wide variety of staffing​ needs, the company has also estimated the number of research scientists required for each development project​ (all cost values are given in millions of​ dollars).

Project Number

Initial Capital​ ($)

Number of Research Scientists

NPV​ ($)

I

10

2

10.1

II

15

3

19.0

III

15

4

22.0

IV

20

3

25.0

V

30

12

60.2

a. Suppose that Orchid has a total capital budget of $60 million. How should it prioritize these​ projects?

b. Suppose that Orchid currently has 12 research scientists and does not anticipate being able to hire more in the near future. How should Orchid prioritize these​ projects?

a. Suppose that Orchid has a total capital budget of $60 million. How should it prioritize these​ projects? The profitability index for Project I is? (Round to two decimal​ places.)
Answer #1

PI=1+NPV/Initial Investment

1.
PI of I=1+10.1/10
=2.01

2.
PI of II=1+19/15
=2.27

3.
PI of III=1+22/15
=2.47

4.
PI of IV=1+25/20
=2.25

5.
PI of V=1+60.2/30
=3.01

6.
Choose Project V,II, III

7.
NPV/Scientist for I=10.1/2
=5.05

8.
NPV/Scientist for II=19/3
=6.33

9.
NPV/Scientist for III=22/4
=5.50

10.
NPV/Scientist for IV=25/3
=8.33

11.
NPV/Scientist for V=60.2/12
=5.02

12.
Choose Project IV, II, III, I